# Deep Emotion Recognition in Textual Conversations: A Survey

### v3

@@ 5.2.1 Metrics for Unbalanced Data
The performance metrics used for Emotion Recognition in Conversations are the Accuracy, Precision, Recall and F-score or F1, the harmonic mean of Precision and Recall. These are commonly used metrics in many classification applications. One can consider the weighted version of these metrics, that take into account the relative frequencies of each class, or micro-averaging, in which all samples equally contribute to the final averaged metric. Since both weighted versions and micro-averaging depend more on the classes with a bigger number of items, those are not good indicators of the performance on the minority classes. Therefore, those should not be used when the classes have the same importance and unequal frequencies. In unweighted macro-averaging the metric is computed for each class, and the results are averaged over all the classes. By maximizing an unweighted metric, one is maximizing the performance of the model in correctly classifying all classes regardless of the number of items they have.
ERC is clearly a task that deals with real-world unbalanced data. As such, a proper evaluation should always be performed on a test set that maintains the real-world unbalance of the data and using unweighted metrics such as macro F-score. Ideally, performance results on each of the individual classes should also be presented. Accuracy, micro-averaging and weighted metrics are mostly irrelevant in what concerns the performance of the models since they are mostly dependent on the majority class, usually neutral. Unweighted metrics that are not affected by dataset unbalance, such as Recall, should also be avoided, unless used in conjunction with those that are, such as Precision.
@@ 5.2.2 Balancing Techniques
Besides the use of apropriate metrics, another way of dealing with dataset unbalance is to apply balancing techniques to the training set. Common such techniques are random under-sampling, removing observations from the majority class and random over-sampling, adding copies to the minority class. More elaborate techniques, such as Smote - Synthetic Minority Oversampling Technique [11], usually improve the results.
Balancing a training set has its advantages, but it is of utmost importance that the test set maintains the original set balance. Otherwise, metrics such as Precision and F-score are artificially improved and the model is unlikely to be effective in an unbalanced real-world dataset. It should be noted that Recall is not affected when training and testing on an artificially balanced dataset so ignoring Precision and looking only at Recall as an evaluation metric will not be a good indicator for real world performance of a model when there is a real-world imbalance among classes.
@@ 5.2.3 Few-Shot Learning
6 Systematic Review Tables of ERC Works
In Tables LABEL:ercmul and 8, a plethora of works on ERC, from 2017 to 2023, are displayed along with the methods they use and their reported performance. Tables are split by datasets and works are split by methods.
@@ 6 Systematic Review Tables of ERC Works
Jiao et al., 2020 [53]* Bi-GRU, Attention 62.7 58.1
Majumder et al., 2019 [78]* Bi-GRU, Attention 62.75
Nwe et al., 2020 [86] Bi-GRU, Attention 63.90
Lu et al., 2020 [75] Bi-GRU, Attention 64.37 60.72
Zhang et al., 2020 [140]* Bi-GRU, Attention 65.3
Zhang and Chai., 2021 [143]* Bi-GRU, Attention 65.37
@@
Partaourides et al., 2020 [89] Bi-LSTM, GRU, Attention 53.0
Poria et al., 2017 [95]* LSTM 54.95
Xing et al., 2022 [134]* Bi-LSTM, GRU, Attention 64.30 60.45
Li et al., 2022a [69] Bi-LSTM 64.59 60.9
Hu et al., 2021 [44] Bi-LSTM, Attention 66.20 58.39
@@
Hazarika et al., 2018 [37]* Memory Network, GRU, Attention 56.13
Hazarika et al., 2018 [36]* Memory Network, GRU, Attention 58.54
Lai et al., 2020 [58]* Memory Network, GRU, Attention 62.43
@@
Ghosal et al., 2019 [31] Graph Neural Network, GRU 64.18 58.10
Ishiwatari et al., 2020 [48] Graph Neural Network, Attention 65.22 60.91 34.42 54.31
Hu et al., 2021 [45]* Graph Neural Network 66.22 58.65
Sheng et al., 2020 [109] Graph Neural Network, Bi-LSTM, Attention 66.61 58.45
Hu et al., 2022 [43]* Graph Neural Network, Bi-GRU 68.18 59.46
Li et al., 2021 [68]* Quantum-inspired Neural Network 59.88 58.00
@@
Hazarika et al., 2021 [38] Transformer, GRU 58.8
Zhong et al., 2019 [149] Transformer, Knowledge Base (KB) 59.56 58.18 34.39 53.37
Zhang et al., 2020 [142] Transformer 61.43 58.97 35.59 54.71
Zhu et al., 2021 [151] Transformer, KB 62.81 68.23 43.12 58.47
Kang and Cho., 2023 [54] Transformer, GNN, Few-Shot Learning 63.16 52.83
Sun et al., 2021 [112] Transformer, Bi-LSTM, GNN 64.10 64.22 36.38
Li et al., 2020 [66] Transformer 64.50 61.94 36.75
Ghosal et al., 2020 [30] Transformer, GRU, KB 65.28 65.21 38.15 58.48
Tu et al., 2023 [116] Transformer, GRU, KB 65.69 66.12 39.47
Shen et al., 2021 [107] Transformer 65.94 62.41 34.73 54.93
Li et al., 2021 [li2021contrast] Transformer, Contrastive Learning [28] 66.18 64.81 39.04 56.29
Zhang et al., 2023 [144] Transformer, Attention 66.35 38.93 61.22
Li et al., 2021 [65] Transformer, GNN, KB 66.96 65.18 38.88 59.75
Xie et al., 2021 [133] Transformer, KB, Attention 66.98 63.24 57.30
Wang et al., 2020 [122] Transformer, Bi-LSTM, CRF 67.10 58.36 63.12
Tu et al., 2023 [117] Transformer, Constrastive Learning 67.16 66.21 40.23 60.96
Zhang et al., 2023 [141] Transformer, GNN 67.68 66.90 40.69 61.84
Shen et al., 2021 [108] Transformer, DAGNN, GRU 68.03 63.65 39.02 59.33
Tu et al., 2023 [118] Transformer, GNN, Contrastive Learning 68.49 65.34 39.20
Yang et al., 2022 [136] Transformer, GNN 68.73 66.18 46.11 59.76
Liu et al., 2022 [74] Transformer, Attention 68.93 65.37
Liang et al., 2022 [72] Transformer, GNN 68.93 64.17 40.05 64.18
Wu et al., 2023 [130]* Transformer, Bi-LSTM, Attention 69.09
Mao et al., 2021 [79]* Transformer 69.23 63.55
Zhang and Li., 2023 [145]* Transformer, GNN 69.6 62.3
Song et al., 2022 [111] Transformer, Contrastive Learning 69.74 67.25 40.94
Yang et al., 2023 [135] Transformer, Adapter, Contrastive Learning 69.81 65.70 38.75 62.51
Zhang et al., 2023 [146]* Transformer - Generative LLM 69.93 71.90 40.05
Hou et al., 2023 [41] Transformer, GNN, Contrastive Learning 70.16 66.65 41.06 62.19
Tu et al., 2023 [115] Transformer, GNN 70.43 66.19 40.51
Li et al., 2023 [64]* Transformer, GNN 71.03 61.77
Lei et al., 2023 [62] Transformer - Generative LLM 71.39 69.15 41.37
Wei et al., 2023 [124]* Transformer, GNN 71.60 63.92
Li et al., 2022b [71]* Transformer, Bi-LSTM 71.77 64.00
Jiang et al., 2023 [51] Transformer, KB 58.66 35.77 54.82
Lee and Choi., 2021 [59] Transformer, GNN 65.36 39.24 61.91
Chen et al., 2023 [12] Transformer, GRU 66.08 39.48
Zhao et al., 2022 [147] Transformer, GRU, Attention 65.37 40.09
Li et al., 2020b [67] Transformer 61.86 35.92
Sun et al., 2021 [112] Transformer, GNN, Bi-LSTM 64.22 36.38
Kim et al., 2020 [55] Transformer 59.20
Table 7: Summary of publications on Emotion Recognition in Conversations (continued)
Table 8: Summary of publications on Emotion Recognition in Conversations, reporting the methods used along with their performance.
Author, Year Methods Dataset, Metric
Transformer Friends, w-F1
Huang et al., 2019 [47] Transformer 80.8
GRU Friends, w-A
Deep et al., 2019 [17] Bi-GRU 75.94
Transformer
Ma et al., 2021 [76] Transformer, Bi-GRU, Hierarchical, Attention 78.2
Jiang et al., 2020 [52] Transformer, Bi-GRU, Hierarchical, Attention 84.1
LSTM EC, m-F1
González et al., 2019 [32] Bi-LSTM, Attention 73.40
Bouchekif et al., 2019 [6] Bi-LSTM, Bi-GRU 74.5
Zhong et al., 2019 [148] Bi-LSTM 74.63
Ge et al., 2019 [29] LSTM, CNN, Attention 75.42
Ma et al., 2019 [77] Bi-LSTM, Attention 75.57
Lee et al., 2019 [61] Bi-GRU 75.81
Ragheb et al., 2019 [102] Bi-LSTM 75.82
Li et al., 2019 [63] Bi-GRU, Bi-LSTM, Attention, XGBoost [13] 75.88
Agrawal et al., 2019 [1] LSTM, GRU, Attention 77.65
Transformer
Xiao et al., 2019 [132] Transformer 76.08
Winata et al., 2019 [128] Transformer, Hierarchical, LSTM 76.77
Huang et al., 2019 [46] Transformer, LSTM, Hierarchical, Attention 77.09
6.1 Benchmark of ERC Works
Several authors resorted to gated neural networks, which can capture dependencies between words and utterances. This effort led to high reported results on the benchmark datasets. Some authors presented works using graph neural networks and the performances obtained are comparable to the ones obtained with gated neural networks. The best performances, however, were yielded when using transformers or transformer-based models, that better capture dependencies in long utterances, combined with the aforementioned state-of-the art deep gated and graph-based network architectures for context modelling. This may reflect the promising role that transformers and pre-trained transformer-based language models play in Natural Language Processing and Emotion Recognition in Conversations and the suitability of recurrent gated and graph-based neural network architectures, as elaborated in previous sections.
7 ERC Challenges Revisited
Practically all works leveraged the information of preceding and sometimes also subsequent [65] utterances, resorting to gated or graph neural networks for context modelling of the utterances that were represented by embeddings from fine-tuned pre-trained language models. Another viable way to perform context modelling is to feed several appended utterances to the pre-trained language model [91]. For speaker specific modelling, also gated [78], graph [31] neural networks and a combination of both [108] architectures were used. Emotion dynamics modelling was less explored. One work considered a GRU to model the emotion dynamics of each party [78]. Some works considered a knowledge base [149] [30] [65] to aid in interpreting the meaning behind commonsense knowledge. However, such knowledge bases are not specific for emotions, which is an extension to consider. The majority of the works resorted to fine-tuning large pre-trained language models to obtain embeddings, which also aids in capturing the meaning behind commonsense knowledge expressions and informal language. Since most benchmark datasets provide a label for each utterance, guiding the models towards classifying emotions per utterance, there are still plenty of research opportunities with regards to real time ERC. The same applies to recognizing emotion causes. Regarding different taxonomies, as benchmarking is made on each dataset separatedely, there are few incentives to match emotions from one dataset to another, which could be a promising practice to consider. Concerning multilingual ERC, the collection and emotion annotation of multilingual datasets is encouraged, and has become more feasible with the advent of multilingual platforms for chat. Interpretability is still a challenge since most works focus on performance rather than explainability.
@@ 8 Conclusion
Research in Emotion Recognition in Conversations is advancing at a high pace and novel application scenarios pose new challenges and opportunities. This survey introduced and elaborated on the challenges pertaining to ERC. Many pipeline processes, from annotation to classification, need to be carefully considered: from leveraging the conversational context, speaker and emotion dynamics modelling, interpreting common sense expressions, informal language and sarcasm, addressing the challenges of real time ERC, recognizing emotion causes, multilingual ERC, different taxonomies to improving interpretability of the models. This survey went on to provide a description of the emotion taxonomies, benchmark datasets, descriptions of the most prominent works in ERC, with explanations of the Deep Learning architectures employed and advisable ERC practices towards better frameworks, elaborating on methods to deal with subjectivity in annotations and modelling and methods to deal with the typically unbalanced ERC datasets. Finally, the survey presented systematic review tables comparing several works regarding the methods used and their performance.
The key takeaways put forward in this survey are now emphasized. Most datasets are unbalanced, motivating the need to leverage techniques and use appropriate performance metrics to address this unbalance. Further exploring mixed emotions with multi label datasets and classifiers is also an aspect to consider. Regarding subjectivity uncertainty in annotations, there are benefits in using inter-annotator (dis)agreement measures to weight or filter annotator’s opinions, to incorporate these measures as additional information to the classifiers and to consider soft label distributions per sample. Concerning deep learning classifier architectures, transformer-based language models play a promising role in Natural Language Processing and Emotion Recognition in Conversations and the most suitable network architectures for the task are gated and graph-based neural networks. ERC in a Generative Framework leveraging LLMs is also promising. Real time ERC, recognizing emotion causes, dealing with different taxonomies across datasets and multilingual ERC remain challenging.
Although research in Emotion Recognition in Conversations has come a long way, there is still much room for improvement, as it can be seen by the performance of the classifiers on the benchmark datasets and the several unexplored directions put forward in this section. While this survey described how current work has addressed several challenges, it also pointed out the opportunities that partly or fully unaddressed challenges constitute.
Finally, important ethical aspects pertaining to Emotion Recognition are now presented. These aspects are, for example and not limited to, whether an Emotion Recognition module should be developed or used for a certain purpose, which data to collect and the subjects behind the data, diversity and inclusiveness, privacy and control, and possible biases and misuses of the application [85]. Research in these directions will benefit the community with better Emotion Recognition in Conversation modules for current and novel applications.

### v4

@@ 5.2.1 Metrics for Unbalanced Data
The performance metrics used for Emotion Recognition in Conversations are Accuracy, Precision, Recall, and F-score or F1, the harmonic mean of Precision and Recall. These are commonly used metrics in many classification applications. One can consider the weighted version of these metrics, that take into account the relative frequencies of each class, or micro-averaging, in which all samples equally contribute to the final averaged metric. Since both weighted versions and micro-averaging depend more on the classes with a bigger number of items, those are not good indicators of the performance of the minority classes. Therefore, those should not be used when the classes have the same importance and unequal frequencies. In unweighted macro-averaging, the metric is computed for each class, and the results are averaged over all the classes. By maximizing an unweighted metric, one is maximizing the performance of the model in correctly classifying all classes regardless of the number of items they have.
ERC is clearly a task that deals with real-world unbalanced data. As such, a proper evaluation should always be performed on a test set that maintains the real-world unbalance of the data and using unweighted metrics such as macro F-score. Ideally, performance results on each of the individual classes should also be presented. Accuracy, micro-averaging, and weighted metrics are mostly irrelevant in what concerns the performance of the models since they are mostly dependent on the majority class, usually neutral. Unweighted metrics that are not affected by dataset unbalance, such as Recall, should also be avoided, unless used in conjunction with those that are, such as Precision.
@@ 5.2.2 Balancing Techniques
Besides the use of appropriate metrics, another way of dealing with dataset unbalance is to apply balancing techniques to the training set. Common such techniques are random under-sampling, removing observations from the majority class, and random over-sampling, adding copies to the minority class. More elaborate techniques, such as Smote - Synthetic Minority Oversampling Technique [8], usually improve the results.
Balancing a training set has its advantages, but it is of utmost importance that the test set maintains the original set balance. Otherwise, metrics such as Precision and F-score are artificially improved and the model is unlikely to be effective in an unbalanced real-world dataset. It should be noted that Recall is not affected when training and testing on an artificially balanced dataset so ignoring Precision and looking only at Recall as an evaluation metric will not be a good indicator for real-world performance of a model when there is a real-world imbalance among classes.
@@ 5.2.3 Few-Shot Learning
6 Systematic Review of ERC Works
In Table LABEL:ercmul, a plethora of works on ERC, since 2017, are displayed along with the methods they use and their reported performance.
@@ 6 Systematic Review of ERC Works
Jiao et al [46]* Bi-GRU, Attention 62.7 58.1
Majumder et al [69]* Bi-GRU, Attention 62.75
Lu et al [67] Bi-GRU, Attention 64.37 60.72
@@
Partaourides et al [79] Bi-LSTM, GRU, Attention 53.0
Poria et al [86]* LSTM 54.95
Xing et al [124]* Bi-LSTM, GRU, Attention 64.30 60.45
Author, Year Methods IEMO MELD ENLP DD
Hu et al [37] Bi-LSTM, Attention 66.20 58.39
@@
Hazarika et al [31]* Memory Network, GRU, Attention 56.13
Hazarika et al [30]* Memory Network, GRU, Attention 58.54
Lai et al [51]* Memory Network, GRU, Attention 62.43
@@
Ghosal et al [26] Graph Neural Network, GRU 64.18 58.10
Ishiwatari et al [41] Graph Neural Network, Attention 65.22 60.91 34.42 54.31
Hu et al [39]* Graph Neural Network 66.22 58.65
Sheng et al [100] Graph Neural Network, Bi-LSTM, Attention 66.61 58.45
Hu et al [36]* Graph Neural Network, Bi-GRU 68.18 59.46
Li et al [60]* Quantum-inspired Neural Network 59.88 58.00
@@
Zhong et al [144] Transformer, Knowledge Base (KB) 59.56 58.18 34.39 53.37
Zhang et al [138] Transformer 61.43 58.97 35.59 54.71
Zhu et al [146] Transformer, KB 62.81 68.23 43.12 58.47
Khang and Cho [47] Transformer, GNN, Few-Shot Learning 63.16 52.83
Sun et al [104] Transformer, Bi-LSTM, GNN 64.10 64.22 36.38
Li et al [59] Transformer 64.50 61.94 36.75
Ghosal et al [25] Transformer, GRU, KB 65.28 65.21 38.15 58.48
Tu et al [108] Transformer, GRU, KB 65.69 66.12 39.47
Shen et al [98] Transformer 65.94 62.41 34.73 54.93
Li et al [61] Transformer, Contrastive Learning [24] 66.18 69.70 49.08 56.46
Zhang et al [139] Transformer, Attention 66.35 38.93 61.22
Li et al [58] Transformer, GNN, KB 66.96 65.18 38.88 59.75
Xie et al [123] Transformer, KB, Attention 66.98 63.24 57.30
Wang et al [114] Transformer, Bi-LSTM, CRF 67.10 58.36 63.12
Tu et al [109] Transformer, Constrastive Learning 67.16 66.21 40.23 60.96
Zhang et al [137] Transformer, GNN 67.68 66.90 40.69 61.84
Shen et al [99] Transformer, DAGNN, GRU 68.03 63.65 39.02 59.33
Yang et al [129] Transformer, GNN, LSTM, KB 68.31 66.25 40.23 60.21
Tu et al [110] Transformer, GNN, Contrastive Learning 68.49 65.34 39.20
Quan et al [91] Transformer, GNN 68.50 63.80 39.19 59.39
Za et al [136] Transformer, KB, GNN, GRU 68.53 63.92 39.56 59.78
Yang et al [128] Transformer, GNN 68.73 66.18 46.11 59.76
Su et al [103] Transformer - Generative LLM, GNN, LSTM 68.90 67.50
Liu et al [66] Transformer, Attention 68.93 65.37
Liang et al [64] Transformer, GNN 68.93 64.17 40.05 64.18
Wu et al [121]* Transformer, Bi-LSTM, Attention 69.09
Duong et al [17] Transformer, GNN 69.10 63.82 39.85
Mao et al [70]* Transformer 69.23 63.55
Zhang and Li [141]* Transformer, GNN 69.6 62.3
Song et al [102] Transformer, Contrastive Learning 69.74 67.25 40.94
Yang et al [126] Transformer, Adapter, Contrastive Learning 69.81 65.70 38.75 62.51
Xu and Yang [125] Transformer, Attention, Contrastive Learning 69.91 67.58 40.79 60.70
Zhang et al [142]* Transformer - Generative LLM 69.93 71.90 40.05
Li et al [57]* Transformer, GNN 70.00 58.94
Hou et al [35] Transformer, GNN, Contrastive Learning 70.16 66.65 41.06 62.19
Yang et al [127] Transformer, Variational Autoencoder [48] 70.22 65.94 62.14
Tu et al [107] Transformer, GNN 70.43 66.19 40.51
Yu et al [132] Transformer, Contrastive Learning 70.41 67.12 40.24
Yun et al [133]* Transformer 70.48 67.37
Li et al [55]* Transformer, GNN 71.03 61.77
Li et al [56]* Transformer 71.04 66.70
Yao and Shi [131]* Transformer, GNN, LSTM, Bi-GRU 71.21 66.25
Lei et al [54] Transformer - Generative LLM 71.39 69.15 41.37
Zhang et al [140]* Transformer, GNN 71.60 62.30
Author, Year Methods IEMO MELD ENLP DD
Wei et al [116]* Transformer, GNN 71.60 63.92
Li et al [63]* Transformer, Bi-LSTM 71.77 64.00
Hu et al [38]* Transformer - Generative LLM, GNN 74.83 68.75
Hong et al [34]* Transformer, Bi-GRU 76.01 40.78
Jiang et al [45] Transformer, KB 58.66 35.77 54.82
Lee and Choi [52] Transformer, GNN 65.36 39.24 61.91
Chen et al [9] Transformer, GRU 66.08 39.48
Zhao et al [143] Transformer, GRU, Attention 65.37 40.09
Sun et al [104] Transformer, GNN, Bi-LSTM 64.22 36.38
Jian et al [44] Transformer, GNN 67.60
Several authors resorted to gated neural networks, which can capture dependencies between words and utterances. This effort led to high reported results on the benchmark datasets. Some authors presented works using graph neural networks and the performances obtained are comparable to the ones obtained with gated neural networks. The best performances, however, were yielded when using transformers or transformer-based models, that better capture dependencies in long utterances, combined with the aforementioned state-of-the-art deep-gated and graph-based network architectures for context modelling. This may reflect the promising role that transformers and pre-trained transformer-based language models play in Natural Language Processing and Emotion Recognition in Conversations and the suitability of recurrent gated and graph-based neural network architectures, as elaborated in previous sections.
6.1 ERC Challenges Revisited
Practically all works leveraged the information of preceding and sometimes also subsequent [58] utterances, resorting to gated or graph neural networks for context modelling of the utterances that were represented by embeddings from fine-tuned pre-trained language models. Another viable way to perform context modelling is to feed several appended utterances to the pre-trained language model [81]. For speaker-specific modelling, also gated [69], graph [26] neural networks and a combination of both [99] architectures were used. Emotion dynamics modelling was less explored. One work considered a GRU to model the emotion dynamics of each party [69]. Some works considered a knowledge base [144, 25, 58, 136] to aid in interpreting the meaning behind commonsense knowledge. The majority of the works resorted to fine-tuning large pre-trained language models to obtain embeddings, which also aids in capturing the meaning behind commonsense knowledge expressions and informal language.
7 Directions for Future Research
We now present suggestions for future work directions.
The first set of directions concerns potentiating the applicability of ERC modules for real-life scenarios. It comprises real-time ERC, recognizing emotion causes, and multilingual ERC, all elaborated in Section 2. There are plenty of research opportunities regarding annotation and modelling efforts.
Further exploring mixed emotions with multi-label datasets and classifiers is also promising. Moreover, training the same model with different datasets by matching emotions from one dataset to another improves the generalization capabilities of the classifiers.
Some surveyed works resorted to knowledge bases. These are, however, not specific to emotions, which would be a useful extension.
Most datasets are unbalanced, motivating the need to leverage techniques and use appropriate performance metrics to address this unbalance, as elaborated in subsection 5.2.
Concerning subjectivity uncertainty in annotations, we highlight using inter-annotator (dis)agreement measures to weight or filter the annotator’s opinions, incorporating these measures as additional information to the classifiers, and considering soft label distributions per sample.
Finally, we encourage further tackling interpretability, as elaborated in Section 2, since most works focus on performance.
@@ 8 Conclusion
Research in Emotion Recognition in Conversations is advancing at a high pace and novel application scenarios pose new challenges and opportunities. Although research in Emotion Recognition in Conversations has come a long way, there is still much room for improvement, as it can be seen by the performance of the classifiers on the benchmark datasets and the several unexplored directions put forward in this section. While we described how current work has addressed several challenges, we also pointed out the opportunities that partly unaddressed challenges constitute.
As main contributions of this survey, we presented partly addressed challenges for ERC, such as recognizing emotion causes, dealing with different taxonomies across datasets, multilingual ERC, and interpretability with associated future work directions. We compiled an extensive list of Deep Learning works in ERC, being the first to simultaneously report their methods, modalities, and performance across various datasets. Finally, our descriptions of Deep Learning methods in the context of ERC provided insights into the suitability of each method for this task, which were not given in previous surveys.
The survey highlights the advantage of leveraging techniques to address unbalanced data, the exploration of mixed emotions, and the benefits of incorporating annotation subjectivity in the learning phase.
Our survey relies on established benchmark datasets. While this aids in comparing different works, it also constitutes a limitation since benchmarking on these datasets may not generalize well to domain-specific applications such as different industries or languages. We also just briefly mention some challenges of real-world applications, not elaborating on the scalability or performance of models in such settings. So our findings may be highly relevant to academic benchmarks but less relevant to domain-specific applications.
Finally, important ethical aspects pertaining to Emotion Recognition are now presented. These aspects are, for example, and not limited to, whether an Emotion Recognition module should be developed or used for a certain purpose, which data to collect and the subjects behind the data, diversity and inclusiveness, privacy and control, and possible biases and misuses of the application [76]. Research in these directions will benefit the community with better Emotion Recognition in Conversation modules for current and novel applications.